RealLifeJuul:

if the length of a rectangle is increased by 30% and the width of the same rectangle is decreased by 30% what is the effect on the area of the rectangle? It is increased by 60% it is unchanged it is decreased by 15% it is decreased by 9%

RealLifeJuul:

@Pixel

RealLifeJuul:

Im thinking its c so far or b

Pixel:

@Shadow

RealLifeJuul:

@Hero

Hero:

@RealLifeJuul for questions like these, it is useful to create a real world scenario, i.e. create the conditions given in the problem "A rectangle of length \(a\)and width \(b\). Then increase \(a\) by \(30\%\) and decrease \( b\) by \(30\%\) as the problem suggests and then assess the area of the rectangle under these conditions compared to the original rectangle.

RealLifeJuul:

oh ok , so if im applying it the way you say to. Then Wouldnt the area of the rectangle remain th same, or am I wrong in assuming that if you increase the length and decrease the width by the same amount technically the rectangle would change at all regarding the area

RealLifeJuul:

??

Hero:

Create the conditions stated in the problem, then assess. Otherwise you're just guessing.

RealLifeJuul:

ok so I made a rectangle with the measurements of 30L and 20W, and went on to multiply those by .30 and added 9 because .30 x 30 = 9 I just added that onto the L and then did the same to the width but this time subtracted and got 14W so my new rectangle taken into account the measurements from before has a new area of 546 and the old area is 600. So now what am I supposed to do here, btw sorry for the explanation being all over the place.

RealLifeJuul:

if I multiply my old area by the answer given and compare it to the new area do you think I can find my answer?

Hero:

To decrease \(a\) by \(30\%\) you compute \(a - 0.30a\).

RealLifeJuul:

I think I got it what I did was 30a x .30 +9 got the L and then I did 20b x .30 -6 and got 39a and 14b I went on to get the area of both subtracted 600 - 546 and then divided 54 over 600 and got a 9% decrease is this right?

Hero:

Nope that's not it either. You're confusing the operations.

Hero:

@RealLifeJuul are you still here?

RealLifeJuul:

yee

Hero:

Suppose the length of a rectangle is \(l = 12\) units and the width is \(w = 6\) units. The area of the rectangle is \(A = 72\) square units If we increase the length by \(30\%\) and decrease the width by \(30\%\) then the length and width of the altered rectangle is \( l = 12 + .30(12) = 15.6\) units and \(w = 6 - .30(6) = 4.2\) units The area of the altered rectangle is \(A = (15.6)(4.2) = 65.52\) square units. Notice that \(0.09(72) = 65.52\). Which means that the Area of the altered rectangle has decreased by \(9\%\) of the original.

RealLifeJuul:

Yeah I see that, so it's not unchanged it's decreased by 9% which I arrived at to with my rectangle.

RealLifeJuul:

or did I do something wrong in my equation?

Hero:

Try posting your work again. I did not follow what you wrote the first time and the operations seemed off.

Hero:

This time focus on typing the operations you computed correctly.

RealLifeJuul:

Ok, so first I got my original rectangle with L being 30 and W being 20 then I multiplied them both by .30 so .30 x 30 = 9 and then did 30+9=39 and got the new Length for the width I did the same thing except I subtracted so .30 x 20 = 6 and then 20-6=14 then I went and got the new area of the altered rectangle that being so LxW 39 x 14 = 546 I also had the area of the unaltered rectangle too. So I have unaltered rectangle A=600 and altered rectangle A= 546. I then went on to subtract the two and then divide with the old rectangles area ontop. so 600-546= 54 and then divided by 54/600 and got .9.

Hero:

Correct.

Hero:

Well, except, you should have gotten \(0.09\) which is \(9\%\)

RealLifeJuul:

Hey thanks for all the help :D

Hero:

You're welcome